If X1 and X2 are independent and identically distributed normal random variables with mean m and variance s2, find the probability distribution function for U=X1-3X2/2.
Answer #1

Answer:-

Given that:-

If X_{1} and X_{2} are independent and identically distributed normal random variables with mean m and variance s2, find the probability distribution function for U=X_{1}-3X_{2}/2

Let X_{1} & X_{2} on independet & identically distribution r.v.s with mean \mu &\sigma ^{2}

X \sim N(\mu,\sigma ^{2})

we know that

M_{X_{1}}(t)=M_{X_{2}}(t)=e^{\mu t+\frac{t^{2}\sigma ^{2}}{2}}

Let U=\frac{X_{1}}{2}-\frac{3X_{2}}{2}

mgf of U

M_{U}(t)=\frac{MX_{1}}{2}-\frac{3X_{2}}{2}(t)

=\frac{MX_{1}}{2}(t),\frac{3X_{2}}{2}(-t)   \begin{Bmatrix} iid \end{Bmatrix}

=MX_{1}(\frac{t}{2}),MX_{2}(\frac{-3t}{2})

=e^{\frac{\mu t}{2}+\frac{t^{2}}{2}\frac{\sigma ^{2}}{4}},e^{\frac{-3\mu t}{2}+\frac{9 t^{2}}{4\times 2}\sigma ^{2}}

=e^{-\mu t+\frac{t^{2}}{2}(\frac{5 }{2}\sigma ^{2})}

Which is mgf  N(-\mu,\frac{5}{2}\sigma ^{2})

Hence y uniduchess theorem

U \sim N(-\mu,\frac{5}{2}\sigma ^{2})